OpenStudy (anonymous):

Find the indicated function values for the following function f(x)5x-2 ----- x-5 a.) f(0) b.) f(1) c.) f(-5) d. f( a+b ) e. why must 5 be excluded from the domain of?

OpenStudy (mertsj):

\[f(x)=\frac{5x-2}{x-5}\]

OpenStudy (mertsj):

Is that what you meant?

OpenStudy (lynfran):

so our function is \[f(x)=\frac{ 5x-2 }{ x-5 }\]and we want to know what f(0).. so where ever we see "x" in the give function we substitute '0' so..\[f(0)=\frac{ 5(0)-2 }{ (0)-5 }\]and we simplify it to get \[f(0)=\frac{ -2 }{ -5 }\]\[f(0)=\frac{ 2 }{ 5 }\]...can u do the rest??

OpenStudy (lynfran):

u just have to follow the same procedure as above

OpenStudy (lynfran):

what u got for f(1) ...??

OpenStudy (lynfran):

\[f(1)=\frac{ 5(1)-2 }{ (1)-5 }\]